# Eagle: Exploring The Design Space for Multimodal LLMs with Mixture of Encoders

### v1

@@ Abstract
Figure 1: Overview of Eagle. Eagle is a family of multimodal large language models (MLLMs) with a mixture of vision encoders. Left: comparisons between Eagle and existing competitive MLLMs with Vicuna-7B [15] and Llama3-8B [3], with Eagle achieving favorable results on all 12 benchmarks. Middle: an evolutionary road map of the design space and advanced training recipes leading to consistent and significant improvements. Right: highlights and core features of Eagle.
@@ 1 Introduction
The success of large language models (LLMs) has triggered significant interest in enabling their visual perception capability, such that they could see, understand, and reason in the real world. At the center of these multimodal large language models (MLLMs) [23] is a typical design where images are converted into a series of visual tokens by the vision encoders and appended with the text embeddings. CLIP [63] is often chosen as the vision encoder since its visual representation is align with the text space by pre-training on image-text pairs. Depending on the architectures, training recipes and the way how vision tokens are injected into the language model, there exist various notable families of MLLMs such as Flamingo [4], BLIP [44, 43, 16], PaLI [11], PaLM-E [18] and LLaVA [54, 52]. Most of these works keep relatively low input resolutions due to the limits on pre-trained vision encoders and LLM sequence length.
@@
Recent studies [47, 53] show that stronger vision encoder design is important for mitigating MLLM hallucinations [50, 77] and improving resolution-sensitive tasks like optical character recognition (OCR). A constellation of works thus focuses on enhancing the capability of the vision encoder. For example, scaling up the pre-training data and parameters of vision encoder [13] or dividing images into low-resolution patches [53, 67]. However, these approaches usually introduce large training resources. An efficient yet powerful strategy is to mixing visual encoders pre-trained with different tasks and input resolutions, either fusing higher resolution encoders with the CLIP encoder [59, 45], sequentially appending features from different encoders [20, 49], or adopting more complex fusion and routing strategies to make the best of different encoders [37, 88]. Such a “mixture-of-vision-experts” mode is shown to be effective. However, a detailed study of its design space with rigorous ablation is still lacking, which motivates us to revisiting this area. Questions such as which vision encoder combination to choose, how to fuse different experts together, and how to adjust training strategies with more vision encoders remain to be answered.
@@
To address the above questions, our work systematically investigates the mixture-of-vision-encoders design space for improved MLLM perception. As shown in Fig. 1, our exploration of the design space consists of the following steps: 1) Benchmarking various vision encoders and searching recipes for higher resolution adaptation; 2) “Apples to apples” comparison between vision encoder fusion strategies; 3) Progressive identification of the optimal combination of multiple vision encoders; 4) Improved vision expert pre-alignment and data mixture. Our study covers the performance of vision encoders pre-trained on different tasks and resolutions (e.g., vision-language alignment [31, 14, 63, 66], self-supervised learning [62], detection [22, 21], segmentation [36], and OCR [38]). We use a round-robin approach to incorporate additional vision experts. Starting with the basic CLIP [63] encoder, we add one additional expert at a time with the best improvement in each round.
@@
  • Unlocking the vision encoders during MLLM training matters. This is in sharp contrast to the LLaVA [54, 52] family and many works that consider multiple vision encoders or teachers [49, 55, 20, 33, 64, 37], where freezing the vision encoders has been a common choice.
  • Some recently proposed fusion strategies [59, 45] do not show significant advantages despite their advanced designs. Instead, we find that straightforward channel concatenation stands out as a simple yet competitive fusion strategy, offering the best efficiency and performance.
  • Incorporating additional vision experts leads to consistent gain, making it a promising path to systematic enhance MLLM perception besides scaling up single encoders. The improvement is particularly pronounced when vision encoders are unlocked.
  • We propose a pre-alignment stage where non-text-aligned vision experts are individually fine-tuned with a frozen LLM before trained together. This stage is found to enhance the MLLM performance significantly under the mixture-of-vision-encoder design.
We finally conclude our findings into a family of MLLMs termed Eagle. Eagle is evaluated on a series of benchmarks, including visual question answering, OCR/document related tasks, and benchmarks tailored for MLLMs. Our model attains state-of-the-art performance across different benchmarks and demonstrates obvious advantages on OCR and document understanding tasks. Using the same pre-train and supervised fine-tuning data from Cambrian-1 [73] - a concurrent family of vision-centric MLLMs sharing similar design spirits, Eagle models overall achieve better performance. We hope that the Eagle can provide a highly performant and easy-to-reproduce MLLM solution to the community.
@@ 2 Design space exploration
In this section, we show how to utilize the advantages of different vision encoders via step-by-step investigations, yielding the Eagle model family. Unlike previous methods focusing on new fusion strategies or architectures among vision encodes, our goal is to identify a set of minimalistic design to fuse different vision encoders supported with detailed ablations, removing any unnecessary parts. As shown in Fig. 2, we start from extending the basic CLIP encoder [63] to a set of vision experts with different architectures, pre-training tasks and resolutions. With these experts, we then compare different fusion architectures and methods, and study how to optimize the pre-training stratgies given more encoders. We also give a detailed analysis in how to select the vision encoders to be integrated. Finally, we put all the findings together and further extend to multiple expert vision encoders with different resolutions and domain knowledge.
@@ 2.1 Base setup
We adopt LLaVA’s [54] model architecture as the basis, which consists of a large language model, a vision encoder, and a projection layer. The projection layer projects the visual embedding from the vision encoder into the text embedding space. We mainly focus on the vision encoder design.
Figure 2: Overview of the Eagle exploration pipeline. We investigate the MLLM design space with multiple vision encoders, aiming to identify optimized design choices and improve the MLLM perception. We include various types of vision experts pre-trained on different tasks and resolutions. We then systematically compare a series of popular fusion strategies under controlled settings, including Sequence Append, Channel Concatenation, LLaVA-HR [59], Mini-Gemini [45], and Deformable Attention [86]. Finally, the optimal fusion design is further extended to multiple vision encoders to form a strong MLLM perception.
Base training data. We use the same pre-training data as LLaVA-$1.51.5$ [52], which consists of 595595595595k image text pairs. For the supervised fine-tuning stage, we collect data from a series of tasks and convert them into multimodal conversations, including: LLaVA-1.5 [52], Laion-GPT4V [1], ShareGPT-4V [8], DocVQA [61], synDog-EN [35], ChartQA [60], DVQA [32], and AI2D [34], resulting in 934k samples.
Implementation details. We first pre-train the model with image-text pairs for one epoch with a batch size of 256, where the whole model is frozen and only the projector layer is updated. In the second stage, we further fine-tune the model on the supervised fine-tuning data for one epoch with a batch size of 128. For this exploration, we employ Vicuna-7B as the underlying language model. The learning rates are set to 1e-3 for the first stage and 2e-5 for the second stage, respectively.
@@ 2.2 Stronger CLIP encoder
We start our exploration with the CLIP model since it has become the main choice for many MLLMs. While CLIP models are known to benefit multimodal tasks, their drawbacks have also been well-studied. For instance, many existing MLLMs [52] tend to use the pre-trained CLIP resolutions (such as $224\times 224224 × 224$ or $336\times 336336 × 336$) as their input resolutions. In these cases, the encoders often fail to capture fine-grained details that are important for resolution-sensitive tasks like OCR and document understanding [47].
@@
To handle increased input resolution, a common practice is to use tiling where input images are divided into tiles and encoded separately [53, 47]. Another simpler method is to directly scale up the input resolution and interpolate the position embeddings of the vision transformer model if needed [9, 10, 6]. We compare these two approaches with frozen/unfrozen vision encoders under different resolutions, with the results shown in Table 1. Our findings can be summarized as follows:
  • Unfreezing the CLIP encoder leads to significant improvement when interpolating to a higher MLLM input resolution that is different from the CLIP pre-training resolution. There is also no performance degradation when resolutions remain the same.
  • When the CLIP encoder is frozen, directly adapting it to a higher MLLM input resolution considerably hurts the performance.
  • Among the compared strategies, directly interpolating to $448\times 448448 × 448$ with an unfrozen CLIP encoder is shown to be both effective and efficient in terms of performance and cost.
@@
  • The best CLIP encoder gets close to InternVL in performance despite the significantly smaller model size (300M vs. 6B) and less pre-training data.
It is worth noting that CLIP-448 allows us to match the setting with LLaVA-HR [59] and InternVL [13] where the CLIP encoders are similarly adapted to take $448\times 448448 × 448$ input and output $10241024$ patch tokens [^1] The latest InternVL 1.0 and InternVL 1.2 models also adopt a pixel shuffle module. We remove it in our experiment to ensure full alignment. See https://github.com/OpenGVLab/InternVL for details.. In the following investigation, we stick to this simple strategy of scaling up the input resolution and unlocking the vision encoder during training.
Table 1: Comparison of different high-resolution adaption methods. Direct interpolation is more efficient and effective. “Res” denotes input image resolution. “#Token/s” denotes the inference speed of the whole framework. “#Token (V)” denotes the number of visual feature tokens fed into LLM. “OCR” denotes OCRBench [57]. “Avg” denotes average. Note that CLIP-448 gets very close to InternVL in terms of performance, despite the significantly smaller model size (300M vs. 6B).
Method Unfreeze  Res #Token/s #Token (V)  GQA  MME MMMU  OCR  SQA POPE  Avg
Original 336 47.2 576 62.46 1488 36.1 369 72.79 86.77 615
Original 336 47.2 576 63.68 1496 36.6 369 72.22 85.9 617
Interpolate 448 47.9 1024 62.45 1484 34.4 285 72.27 86.54 598
Interpolate 448 47.9 1024 64.62 1531 35.7 496 72.77 87.88 645
Interpolate 672 44.5 2304 64.51 1492 36.2 492 72.86 87.49 641
Tiled-input 672 44.8 2304 63.59 1455 34.8 445 71.94 87.83 625
InternVL [13] 448 46.2 1024 65.12 1521 36.2 527 72.46 87.13 649
Table 2: Detailed setting of the pre-trained vision experts.
Expert Category Resolution Token Number Post-Processing Variant Hyper-links
CLIP Image-Text Matching 448 1024 None VIT-L
ConvNeXt Image Classification 1024 1024 None ConvNeXt-XXL
EVA-02 Object Detection 1024 1024 Resize EVA-02-Large
Pix2Struct Text Recognition 1024 1024 Resize Pix2Struct-02-Large
DINOv2 Self-supervised 448 1024 None dinov2_vitl14_reg
SAM Image Segmentation 1024 1024 Pixel-unshuffle SAM-Large
@@ 2.3 Vision experts
To better establish the foundation for multi-vision expert fusion, we extend the toolbox with vision experts pre-trained on different tasks and resolutions, and verify our findings on high-resolution adaptation with these experts. This also helps us identify the distinct advantages of different experts. We collect a set of vision experts, including: (1) Vision-Language Alignment: CLIP [63] and ConvNeXt [76] from OpenCLIP [31, 66]. (2) Object-Centric: EVA-02 [22, 21] pre-trained on detection datasets. (3) OCR: Pix2Struct [38].(4) Segmentation: SAM [36]. (5) Self-supervised: DINOv2 [62]. The detailed input resolution and checkpoint of each vision encoder can be found in Table 2. We resize the output 2D feature maps of each vision encoder using bilinear interpolation and pixel shuffle [68] to ensure that the visual token number equals $10241024$.
The results in Table 3 reveal that MLLMs with these task-specific vision encoders achieve optimal performance in their pre-training domains. EVA-02 excels in the object hallucination evaluation benchmark POPE [46] and general visual question answering benchmark GQA [30]. CLIP and ConvNeXt perform well across all benchmarks, benefiting from their training on large-scale image-text pairs using contrastive loss. Conversely, while Pix2Struct excels in text recognition, it shows limited capability in object recognition and general VQA tasks, like POPE and GQA. DINOv2 and SAM, pre-trained with self-supervised learning and semantic segmentation respectively, struggle with text recognition tasks. Note that the conclusion made in Sec. 2.2 aligns with the results here. Both ConvNeXt and EVA-02 show significant performance gains when the vision encoders are updated during training.
Table 3: Comparison between different vision experts as the MLLM encoders. “Res” denotes resolution. “OCR” denotes OCRBench [57].
Category Vision Tower Unfreeze    Res  GQA  MME MMMU  OCR  SQA POPE TextVQA
VL Alignment ConvNeXt 1024 63.71 1473 34.2 402 71.92 87.42 66.95
1024 63.71 1433 37.0 527 72.01 87.91 71.66
Segmentation SAM 1024 54.87 1193 34.8 32 70.34 83.31 45.04
1024 60.05 1305 35.9 38 70.64 86.62 45.71
Object Detection EVA-02 1024 63.54 1422 35.8 58 70.86 88.50 47.86
1024 64.93 1474 35.7 387 71.78 88.96 59.79
Text Recognition Pix2Struct 1024 52.47 1217 35.7 443 69.61 77.54 56.56
1024 54.05 1220 34.3 425 70.29 79.21 57.05
Self-supervised DINOv2 448 60.81 1269 34.0 39 70.76 86.42 45.71
448 64.06 1371 34.2 37 71.35 87.29 46.77

### v2

# Eagle: Exploring The Design Space for Multi- modal LLMs with Mixture of Encoders

@@ Abstract
Figure 1: Overview of Eagle. Eagle is a family of multimodal large language models (MLLMs) with a mixture of vision encoders. Left: comparisons between Eagle and existing competitive MLLMs with Vicuna-13B (Chiang et al., 2023), with Eagle achieving favorable results on all 13 benchmarks. Middle: an evolutionary road map of the design space and advanced training recipes leading to consistent and significant improvements. Right: highlights and core features of Eagle.
@@ 1 Introduction
The success of large language models (LLMs) has triggered significant interest in enabling their visual perception capability, such that they could see, understand, and reason in the real world. At the center of these multimodal large language models (MLLMs) (Fei et al., 2024) is a typical design where images are converted into a series of visual tokens by the vision encoders and appended with the text embeddings. CLIP (Radford et al., 2021) is often chosen as the vision encoder since its visual representation is aligned with the text space by pre-training on image-text pairs. Depending on the architectures, training recipes, and the way how vision tokens are injected into the language model, there exist various notable families of MLLMs such as Flamingo (Alayrac et al., 2022), BLIP (Li et al., 2022; 2023d; Dai et al., 2024), PaLI (Chen et al., 2023e), PaLM-E (Driess et al., 2023) and LLaVA (Liu et al., 2023d; c). Most of these works keep relatively low input resolutions due to the limits on pre-trained vision encoders and LLM sequence length.
@@
Recent studies (Li et al., 2024c; Liu et al., 2024a) show that stronger vision encoder design is important for mitigating MLLM hallucinations (Liu et al., 2023a; Wu et al., 2024) and improving resolution-sensitive tasks like optical character recognition (OCR). A constellation of works thus focuses on enhancing the capability of the vision encoder. For example, scaling up the pre-training data and parameters of vision encoder (Chen et al., 2023f) or dividing images into low-resolution patches (Liu et al., 2024a; Shi et al., 2024). However, these approaches usually introduce large training resources. An efficient yet powerful strategy is to mix visual encoders pre-trained with different tasks and input resolutions, either fusing higher resolution encoders with the CLIP encoder (Luo et al., 2024; Li et al., 2024b), sequentially appending features from different encoders (Fan et al., 2024; Lin et al., 2023b; Karamcheti et al., 2024; Tong et al., 2024), or adopting more complex fusion and routing strategies to make the best of different encoders (Lee et al., 2024; Zong et al., 2024). In addition, Prismatic VLM (Liu et al., 2024b) incorporates multiple vision encoders with channel concatenation as part of their design space exploration. These “mixture-of-vision-experts” strategies are shown to be effective. However, a detailed study focusing on their designs is still lacking.
@@
To address the above questions, our work systematically investigates the mixture-of-vision-encoders design space for improved MLLM perception. As shown in Fig. 1, our exploration of the design space consists of the following steps: 1) Benchmarking various vision encoders and searching recipes for higher resolution adaptation; 2) “Apples to apples” comparison between vision encoder fusion strategies; 3) Progressive identification of the optimal combination of multiple vision encoders; 4) Improved vision expert pre-alignment and data mixture. Our study covers the performance of vision encoders pre-trained on different tasks and resolutions (e.g., vision-language alignment (Ilharco et al., 2021; Cherti et al., 2023; Radford et al., 2021; Schuhmann et al., 2022), self-supervised learning (Oquab et al., 2023), detection (Fang et al., 2023b; a), segmentation (Kirillov et al., 2023), and OCR (Lee et al., 2023)). We use a round-robin approach to incorporate additional vision experts. Starting with the basic CLIP (Radford et al., 2021) encoder, we add one additional expert each time with the best improvement in each round.
@@
  • Unlocking the vision encoders during MLLM training matters. This is in sharp contrast to the LLaVA (Liu et al., 2023d; c) family and many works that consider multiple vision encoders or teachers (Lin et al., 2023b; Liu et al., 2024b; Fan et al., 2024; Kar et al., 2024; Ranzinger et al., 2024; Lee et al., 2024), where freezing the vision encoders has been a common choice.
  • Some recently proposed fusion strategies (Luo et al., 2024; Li et al., 2024b) do not show significant advantages despite their advanced designs. Instead, we find that straightforward channel concatenation stands out as a simple yet competitive fusion strategy, offering the best efficiency and performance.
  • Incorporating additional vision experts leads to consistent gain, making it a promising path to systematically enhance MLLM perception besides scaling up vision encoders. The improvement is particularly pronounced when vision encoders are unlocked.
  • We propose a pre-alignment stage where non-text-aligned vision experts are individually fine-tuned with a frozen LLM before being trained together. This stage is found to enhance the MLLM performance significantly under the mixture-of-vision-encoder design.
We finally conclude our findings into a family of MLLMs termed Eagle. Eagle is evaluated on a series of benchmarks, including visual question answering, OCR/document-related tasks, and benchmarks tailored for MLLMs. Our model attains state-of-the-art performance across different benchmarks and demonstrates obvious advantages on OCR and document understanding tasks. Using the same pre-train and supervised fine-tuning data from Cambrian-1 (Tong et al., 2024) - a concurrent family of vision-centric MLLMs sharing similar design spirits, Eagle models overall achieve better performance. We hope that the Eagle can provide a highly performant and easy-to-reproduce MLLM solution to the community.
@@ 2 Design space exploration
In this section, we show how to utilize the advantages of different vision encoders via step-by-step investigations, yielding the Eagle model family. Unlike previous methods focusing on new fusion strategies or architectures among vision encodes, our goal is to identify a set of minimalistic design to fuse different vision encoders supported with detailed ablations, removing any unnecessary parts. As shown in Fig. 2, we start by extending the basic CLIP encoder (Radford et al., 2021) to a set of vision experts with different architectures, pre-training tasks, and resolutions. With these experts, we then compare different fusion architectures and methods and study how to optimize the pre-training strategies given more encoders. We also give a detailed analysis of how to select the vision encoders to be integrated. Finally, we put all the findings together and further extend to multiple expert vision encoders with different resolutions and domain knowledge.
@@ 2.1 Base setup
We adopt LLaVA-1.5’s (Liu et al., 2023c) model architecture as the basis, which consists of a large language model (Vicuna-v1.5 7B (Chiang et al., 2023)), a vision encoder, and a projection layer. The projection layer projects the visual embedding from the vision encoder into the text embedding space.
Base training data. We adopt the same pre-training data (LLaVA-595k) as LLaVA-1.5 (Liu et al., 2023c) for the first pre-training stage, which consists of 595595595595k image text pairs. To fully examine the potential of different vision experts and fusion methods, instead of using the SFT data from LLaVA-1.5 (Liu et al., 2023c), we collect data from a series of tasks and convert them into multimodal conversations for the supervised fine-tuning (SFT) stage, denoted as Eagle1.8M in Table 1.
Table 1: Composition of the base supervised fine-tuning data (Eagle1.8M).
Total Data Size Data Source
1,809k LLaVA-1.5 (665k) (Liu et al., 2023c), DocVQA (39k) (Mathew et al., 2021), synDog-EN (50k) (Kim et al., 2022), ChartQA (28k) (Masry et al., 2022), DVQA (25k) (Kafle et al., 2018), AI2D (15k) (Kembhavi et al., 2016a), ShareGPT-4V (100k) (Chen et al., 2023b), laion-GPT4V (11k) (lai, 2023), LVIS-Instruct4V (220k) (Wang et al., 2023a), LRV-Instruct (150k) (Liu et al., 2023b), Geo170k (120k) (Gao et al., 2023), LLaVAR (20k) (Zhang et al., 2023), Visual7W (70k) (Zhu et al., 2016), Open-Hermes 2.5 (300k) (Teknium, 2023)
Base training recipes. We start from the LLaVA-1.5 (Liu et al., 2023c) recipe where the model is first pre-trained with image-text pairs for one epoch with a batch size of 256. The whole model is frozen and only the projector layer is updated in this pre-training stage. In the second stage, we further fine-tune the model on the multi-modal conversation data for one epoch with a batch size of 128. The learning rates are set to be 1e-3 for the first stage and 2e-5 for the second stage, respectively.
Base evaluation. To conduct a comprehensive comparison of various methods, we adopt 11 distinct benchmarks that span multiple tasks. These benchmarks include 1) General VQA tasks: GQA (Hudson & Manning, 2019), VizWiz (Gurari et al., 2018), MME (Fu et al., 2023), SEED (Li et al., 2023c); 2) OCR/document/chart understanding: OCRBench (Liu et al., 2023f), DocVQA (Mathew et al., 2021), ChartQA (Masry et al., 2022); 3) vision-centric tasks: POPE (Li et al., 2023e), RealWorldQA (xAI, 2024); 4) knowledge-based tasks: ScienceQA (Saikh et al., 2022), AI2D (Kembhavi et al., 2016b). To obtain an average score, we normalize each benchmark to a total score of 1,000 and then calculate the average score across all benchmarks.
Figure 2: Overview of the Eagle exploration. In this work, we explore the design space of Multi-Modal Large Language Models (MLLMs) with multiple vision encoders, aiming to identify optimized design choices and enhance MLLM perception. We collect a range of vision experts and adapt them for integration into MLLMs. A systematic comparison of popular fusion paradigms is then conducted under controlled settings. After identifying discrepancies between vision experts pre-trained on different tasks, we optimize the pre-training strategy through a pre-alignment stage and use round-robin searching to determine the optimal combination of vision encoders.
@@ 2.2 Stronger CLIP encoder
We start our exploration by upgrading the vanilla CLIP Radford et al. (2021) model since it has become a standard choice for most of the MLLMs Liu et al. (2023d; c). While CLIP models are known to benefit multimodal tasks via the text-image alignment, they also have inherent drawbacks. For instance, many existing MLLMs (Liu et al., 2023c) tend to use the pre-trained CLIP resolutions (such as $224\times 224224 × 224$ or $336\times 336336 × 336$) as their input resolutions. In these cases, the encoders often fail to capture fine-grained details that are important for resolution-sensitive tasks like OCR and document understanding (Li et al., 2024c).
@@
To handle increased input resolution, a common practice is to use tiling where input images are divided into tiles and encoded separately (Liu et al., 2024a; Li et al., 2024c), or just interpolate the position embedding of the vision transformer model to fit high-resolution inputs (Chen et al., 2023c; d; Beyer et al., 2024). We compare these two approaches with frozen/unfrozen vision encoders under different resolutions, with the results shown in Table 2. Our findings can be summarized as follows:
  • Updating the CLIP encoder during SFT significantly improves performance at higher resolutions but slightly reduces it when using the pre-training resolution.
  • Interpolating CLIP encoder to fit the input size of $448\times 448448 × 448$ offers a strong balance between efficiency and performance, trailing the $672\times 672672 × 672$ version with less than half the tokens.
@@
  • Despite its smaller size (0.3B vs. 5.9B) and less pre-training data, the CLIP encoder gets close with interpolation approaches InternVL’s (Chen et al., 2023f) performance under the same setting.
Table 2: Comparing different high-resolution adaption methods. “#Tok(V)” denotes the number of visual tokens. “#Parames”, “FLOPs” and “Img/Sec” denote the model size, complexity and throughput (bs=4) of the vision encoder.
Method Unfreeze Res. #Tok(V) #Params FLOPs Img/Sec Avg.
Original 336 576 0.3B 119G 197.2 616.5
Original 336 576 0.3B 119G 197.2 562.6
Interpolate 448 1024 0.3B 214G 119.5 589.7
Interpolate 448 1024 0.3B 214G 119.5 670.5
Interpolate 672 2304 0.3B 480G 56.3 674.2
Tiled-input 672 2304 0.3B 476G 51.6 673.9
InternVL 448 1024 5.9B 5669G 13.52 661.9
InternVL 448 1024 5.9B 5669G 13.52 671.5
Based on the results, we can see that Direct interpolation to $448\times 448448 × 448$ can achieve competitive performance while being more efficient. We thus use the CLIP encoder with a $448\times 448448 × 448$ input resolution while unlocking the encoder during the SFT stage.
@@ 2.3 Vision experts
To better establish the foundation for multi-vision expert fusion, we extend the toolbox with vision experts pre-trained on different tasks and resolutions, and verify our findings on high-resolution adaptation with these experts. This also helps us identify the distinct advantages of different experts. We collect a set of vision experts, including: (1) Vision-Language Alignment: CLIP (Radford et al., 2021) and ConvNeXt (Liu et al., 2022) from OpenCLIP (Ilharco et al., 2021; Schuhmann et al., 2022). (2) Object-Centric: EVA-02 (Fang et al., 2023b; a) pre-trained on detection datasets. (3) OCR: Pix2Struct (Lee et al., 2023).(4) Segmentation: SAM (Kirillov et al., 2023). (5) Self-supervised: DINOv2 (Oquab et al., 2023). We resize the output 2D feature maps of each vision encoder using bilinear interpolation or pixel shuffle (Shi et al., 2016) to ensure that the visual token number equals $10241024$.
Table 3: Comparison between different vision experts as the MLLM encoders.
Category Vision Encoder   Res. Post-process Unfreeze   Avg. Model Link
VL Alignment ConvNeXt 1024 None 654.6 ConvNeXt-XXL
682.1
Segmentation SAM 1024 Pixel Unshuffle 486.2 SAM-ViT-Large
510.5
Object Detection EVA-02 1024 Resize 543.7 EVA-02-L-Det
639.1
Text Recognition Pix2Struct 1024 Resize 598.6 Pix2Struct-02-Large
606.2
Self-Supervised DINOv2 448 None 520.7 ViT-L/14-Reg
537.3
Results in Table 3 show that unfreezing the vision experts again leads to consistent improvement, which is aligned with Sec. 2.2. In addition, results in Table 10 (see Appendix A.1) further demonstrate that MLLMs with these task-specific vision encoders achieve optimal performance in their pre-training domains. EVA-02 excels in the object hallucination evaluation benchmark POPE and general visual question answering benchmark GQA. CLIP and ConvNeXt perform well across all benchmarks, benefiting from their training on large-scale image-text pairs using contrastive loss. Conversely, while Pix2Struct excels in text recognition, it shows limited capability in object recognition and general VQA tasks, like POPE and GQA. DINOv2 and SAM, pre-trained with self-supervised learning and segmentation, struggle with text recognition tasks.